Evaluate the definite integral:
∫−4−7(x−1+3x)dx=
1
Expert's answer
2020-12-09T08:05:30-0500

"\int_{-4}^{-7} (x - 1 + 3x)dx\\
Solution:\\
\left[\LARGE{\frac{x^{1+1}}{1 + 1} - 1x + 3\frac{x^{1+1}}{1 +1}}\right]_{-4}^{-7}\\
\left[\LARGE{\frac{x^{2}}{2} - 1x + 3\frac{x^{2}}{2}}\right]_{-4}^{-7}\\
\left[\LARGE{\frac{x^{2}}{2} + 3\frac{x^{2}}{2} -x}\right]_{-4}^{-7}\\
\left[\LARGE{4\frac{x^{2}}{2} -x}\right]_{-4}^{-7}\\
\left[\LARGE{\frac{4x^2 - 2x}{2}}\right]_{-4}^{-7}\\
\left(\LARGE{\frac{4(-7)^2 - 2(-7)}{2}}\right) - \left(\LARGE{\frac{4(-4)^2 - 2(-4)}{2}}\right)\\
105 - 36\\
= 69"